If𝑓(π‘₯,𝑦)=10π‘₯5βˆ’3π‘₯2𝑦6βˆ’4𝑦4.𝐹𝑖𝑛𝑑𝑓 (π‘₯,𝑦),𝑓 (π‘₯,𝑦), π‘₯π‘₯ π‘₯𝑦 𝑓 (1, βˆ’2), 𝑓 (1, 2)


1
Expert's answer
2021-04-29T17:13:50-0400

"\ud835\udc53(\ud835\udc65,\ud835\udc66)=10\ud835\udc65^5\u22123\ud835\udc65^2\ud835\udc66^6\u22124\ud835\udc66^4"


Differentiate w.r.t x

"f(x,y)_x=50x^4=6xy^6"


Again differentiate w.r.t x-


"f(x,y)_{xx}=200x^3-6y^6"


"f(1,2)_{xx}=200(1)^3-6(2)^6=200-384=-184"


"f(1,-2)_{xx}=200(1)^3-6(-2)^6=200-384=-184"


Now differentiate w.r.t y-


"f(x,y)_y=-18x^2y^5-16y^3"


Again differentiate w.r.t. y-


"f(x,y)_{yy}=-90x^2y^4-48y^2"


"f(1,2)_{yy}=-90(1)^2(2)^4-48(2)^2=-1440-192=-1632"


"f(1,-2)_{yy}=-90(1)^2(-2)^4-48(-2)^2=-1140-192=-1632"